# Automating the Refinement of Reinforcement Learning Specifications

### v1

@@ 1 Introduction
Reinforcement Learning (RL) algorithms have made tremendous strides in recent years sutton2018reinforcement; SilverHMGSDSAPL16; MnihKSRVBGRFOPB15; LevineFDA16. However, most algorithms assume access to a scalar reward function that must be carefully engineered to make environments amenable to RL—a practice known as reward engineering rewardengineering. This creates challenges in applying RL to new environments where useful reward functions are hard to construct. Furthermore, scalar Markovian rewards cannot provide sufficient feedback for certain tasks abel2021reward; bowling2023settling, leading to growing interest in non-Markovian reward functions DBLP:journals/corr/LiVB16; dirl; alur2023specification.
To make non-Markovian rewards tractable, it is standard to represent them via logical specification formulas that capture the intended task. These approaches, known as specification-guided reinforcement learning AksarayJKSB16; LiVB17; IcarteKVM18; spectrl; dirl, derive reward functions from logical specifications. However, this creates two challenges: (i) providing specifications granular enough to guide RL algorithms, and (ii) defining accurate labeling functions mapping environment states to specification predicates. Users often create coarse specifications or labeling functions that, while logically correct, provide insufficient guidance for learning.
@@
AutoSpec works with SpectRL specifications; boolean and sequential combinations of reach-avoid tasks spectrl. Any SpectRL specification decomposes into an abstract graph where edges specify reach-avoid tasks dirl. AutoSpec identifies problematic edges and applies targeted refinements: either modifying the labeling function for regions (Figure 1) or restructuring the graph to add alternative paths. These problematic edges are identified by employing an exploration-guided strategy that utilizes empirical trajectory data to identify edges in the abstract graph whose reach-avoid tasks (i.e. initial, target or unsafe regions) make it hard to learn a good RL policy. For instance, in Figure  1, the initial MID region of the MID-GOAL edge in the abstract graph is under-specified, as it overlaps with a trap state. The trap state is not immediately obvious as there is a path from MID to GOAL. By analyzing explored traces, AutoSpec identifies problematic start states in MID and refines the region to exclude the trap (as shown in Figure 1), thereby refining the logical reach-avoid specification associated to the MID-GOAL edge.
We prove that all refinements maintain soundness, where satisfaction of the refined specification implies satisfaction of the original. AutoSpec integrates with existing SpectRL-compatible algorithms as demonstrated with DiRL dirl and LSTS lsts.
@@
Related work. Recent years have seen substantial progress in solving RL tasks specified via logical specifications AksarayJKSB16; LiVB17; IcarteKVM18; CamachoIKVM19; GiacomoIFP19; HasanbeigKA22; HasanbeigKAKPL19; HahnPSSTW19; spectrl; dirl; 0005T19. Many works consider different fragments of Linear Temporal Logic (LTL) or their variants for specifying RL tasks. IcarteKVM18; CamachoIKVM19 consider tasks that can be specified using deterministic finite automata (DFA) and solve them by reward machines, which decompose these tasks and translate them into a reward function. The reward function can then be used to train existing RL algorithms. LiVB17 considers a variant of LTL called TLTL for specifying tasks and propose a method for translating these specifications into continuous reward functions. HasanbeigKA22; HasanbeigKAKPL19; HahnPSSTW19 study the translation of tasks specified in LTL into reward functions. alur2022framework examines the theoretical questions related to the translation of logical specifications into reward functions.
@@
spectrl defines the specification language SpectRL, a finitary fragment of LTL and provides justification for using this language to define specifications for RL tasks. A compositional method that decomposes SpectRL specifications into an abstract graph and constructs a reward function for each abstract graph edge was proposed in dirl. The approach by perfectrm focuses on discovering optimal reward structures through environmental exploration and reward analysis. Compositional methods are further explored by subtask-removal, who propose removing unfulfillable subtasks, and neary2023verifiable, who introduce verification techniques to certify learned policies. vzikelic2024compositional propose Claps, a compositional method for learning neural network policies with formal guarantees on the satisfaction of SpectRL specifications, thus advancing the applicability to safety-critical RL applications by utilizing prior methods for learning reach-avoid policies with formal guarantees LechnerZCH22; ZikelicLHC23; ChatterjeeHLZ23.
@@
Recent advancements also include LTL2Action ltl2action, which translates LTL specifications into sequences of tasks for RL agents. Other recent approaches, such as  qiu2023instructing and DeepLTL jackermeier2025deepltl, leverage goal-conditioned RL and automata-based architectures to solve complex LTL and $\omega$-regular tasks zero-shot or in multi-task settings. Trainify trainify employs counterexample-guided abstraction and refinement (CEGAR) to iteratively improve policies by addressing failure cases identified through counterexamples. However, these works primarily focus on learning policies for fixed, well-defined specifications. In contrast, AutoSpec studies the problem of automated logical specification refinement towards improving reward functions obtained by translation from coarse logical RL specifications. Thus, our work is complementary to the works on RL from logical specifications and can be integrated into off-the-shelf specification-guided RL algorithms to improve the performance of learned agents.
@@ 2 Preliminaries
MDPs. A Markov Decision Process (MDP) is a tuple $M=(S,A,P,R,\gamma)$, where $S\subseteq\mathbb{R}^{n}$ is the state space, $A\subseteq\mathbb{R}^{m}$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probabilistic transition function, $R$ is the (possibly non-Markovian) reward function, and $\gamma$ is the discount factor. Let $\eta:S\rightarrow[0,1]$ be the initial state distribution. A trajectory $\zeta$ in $M$ is a sequence of states and actions $\zeta=s_{0},a_{0},s_{1},a_{1},\ldots$ where $s_{i}\in S$ and $a_{i}\in A$. We use $\mathcal{Z}$ to denote the set of all trajectories in $M$ and $\mathcal{Z}_{f}$ to denote the set of all finite trajectories in $M$, which are finite prefixes of trajectories ending in states. A (pure) policy $\pi:\mathcal{Z}_{f}\rightarrow A$ assigns an action to each finite trajectory, and a non-Markovian reward $R:\mathcal{Z}_{f}\rightarrow\mathbb{R}$ assigns a reward to a finite trajectory. The MDP $M$ under any policy $\pi$ gives rise to a probability space over the set of all trajectories in the MDP Puterman94. We use $\mathbb{P}^{\pi}$ and $\mathbb{E}^{\pi}$ to denote the probability measure and the expectation operator in this probability space, respectively.
Logical specifications for Reinforcement Learning. In this work, we are solving RL tasks defined by logical specifications. Formally, a logical specification (or, simply, a specification) is a boolean function $\phi:\mathcal{Z}\rightarrow\{\textrm{true},\textrm{false}\}$ which specifies whether a trajectory in the MDP satisfies the specification. We write $\zeta\models\phi$ whenever a trajectory $\zeta$ satisfies the specification $\phi$. The objective of a specification-guided RL task is to find a policy $\pi^{*}$ that maximizes the probability of satisfying the given specification $\phi$, i.e. $\pi^{*}\in\textit{argmax}_{\pi}\mathbb{P}^{\pi}[\zeta\models\phi]$. Specification-guided RL algorithms use the specification to create a dense reward that guides the policy search, and therefore outperform algorithms that cannot leverage the specification for learning and instead require manual reward engineering dirl; lsts.
SpectRL specification logic. We consider RL tasks specified in the SpectRL specification logic. SpectRL spectrl is a fragment of Linear Temporal Logic (LTL) which consists of all boolean and sequential combinations of reach-avoid tasks. Formally, a specification in SpectRL is defined in terms of predicates and specification formulas. An atomic predicate is a function $a:S\rightarrow\{\textrm{true},\textrm{false}\}$ which defines a set of states that satisfy the atomic predicate. A predicate is a boolean combination of atomic predicates, i.e. $b:=a\,\mid\,b_{1}\land b_{2}\,\mid\,b_{1}\lor b_{2}$, where $a$ is an atomic predicate and $b_{1}$ and $b_{2}$ are predicates. Specification formulas in SpectRL are defined by the grammar
@@
where $b$ is a predicate and $\phi_{1}$ and $\phi_{2}$ are specification formulas. Intuitively, $\textrm{achieve }b$” requires the agent to reach a state in which the predicate $b$ is satisfied. The clause $\phi_{1}\textrm{ ensuring }b$” requires the agent to satisfy the specification $\phi$ while only visiting states in which the predicate $b$ is satisfied. The clause $\phi_{1};\phi_{2}$” requires the agent to first satisfy specification $\phi_{1}$ and then satisfy specification $\phi_{2}$. The clause $\phi_{1}\textrm{ or }\phi_{2}$” requires satisfaction of at least one of $\phi_{1}$ or $\phi_{2}$. See spectrl for the formal definition of the semantics of each clause.
Abstract graphs for SpectRL specifications. It was shown in dirl that each specification written in the SpectRL specification logic can be translated into an equivalent abstract graph. An abstract graph is a directed acyclic graph (DAG) whose vertices represent sets of MDP states and whose edges are annotated with sets of safe MDP states. Hence, each abstract graph edge defines a reach-avoid specification, where the task is to reach the set of states defined by the target vertex of the edge starting from the set of states defined by the source vertex of the edge, while staying within the set of safe states defined by the edge.
@@
Given a trajectory $\zeta$ in the MDP and an abstract graph $G=(V,E,\beta,s,t)$, we say that $\zeta$ satisfies abstract reachability for $G$ (written $\zeta\models G$) if it gives rise to a path in $G$ that traverses $G$ from $s$ to $t$ and satisfies the reach-avoid specifications of every traversed edge. It was shown in dirl that, given any SpectRL specification $\phi$, one can construct an abstract graph $G$ such that $\zeta\models\phi$ if and only if $\zeta\models G$ holds for each trajectory $\zeta$ in the MDP. Hence, solving an RL task for a SpectRL specification reduces to solving an abstract reachability task in the abstract graph $G$.
Problem statement. Given an MDP $M$ and a SpectRL specification $\phi$, our goal is to learn a policy $\pi$ such that the probability $\mathbb{P}^{\pi}[\zeta\models\phi]$ of a trajectory induced by the policy satisfying the specification is maximized.
@@ 3 Automated Refinement of RL Specifications
Overview of AutoSpec.
Algorithm 1 shows the complete AutoSpec framework. The algorithm takes as input an MDP $M$, a SpectRL specification $\phi$, a satisfaction threshold $p$, and any specification-guided RL algorithm $\mathcal{A}$. It first translates $\phi$ into an abstract graph $G$ and uses $\mathcal{A}$ to learn policies for the graph edges. For each edge $e$ where $\mathcal{A}$ learned a policy but that policy fails to achieve satisfaction probability $p$, AutoSpec applies four refinement procedures in sequence: SeqRefine, AddRefine, PastRefine, and OrRefine. This ordering reflects increasing levels of structural modification; from local predicate adjustments to graph topology changes. The first refinement that successfully improves performance above threshold $p$ is applied, the graph is updated, and policies are relearned before proceeding to the next edge.
@@
AutoSpec iterates through all edges $e=u\rightarrow u^{\prime}$ of the abstract graph $G$ for which the specification-guided RL algorithm $\mathcal{A}$ has learned a policy but for which the reach-avoid task satisfaction probability is below the provided probability threshold pp. For each such edge, AutoSpec performs four refinement procedures that focus on different possible reasons for the edge $e=u\rightarrow u^{\prime}$ being challenging for learning a satisfactory policy.
@@ 3.1 Specification Refinement Subprocedures
ReachRefine collects all states along sampled trajectories that successfully reached the goal region $b$. The refined goal region is computed as $b_{r}=b\cap\text{ConvexHull}(\text{reached states})$, effectively excluding unreachable portions of the original target region.
AvoidRefine collects states where trajectories entered unsafe regions (complement of $c$). The refined safe region is computed as $c_{r}=c\setminus\text{ConvexHull}(\text{last }k\text{ unsafe states})$, where $k$ is a hyperparameter controlling how much of the trajectory tail to consider. This removes demonstrated unsafe areas from the safety region. We cannot relax or reduce the avoid regions during refinement because it would break soundness. In the relaxation, a policy can enter states that were part of the avoid in the original specification. Thus, the refined specification in that case is not a subset of the original specification.
@@
OrRefine: Exploiting Alternative Paths. The fourth refinement addresses blocked or infeasible direct paths by leveraging the existing graph structure. When the path through edge $e=u\rightarrow u^{\prime}$ cannot be made satisfactory, OrRefine identifies alternative parents of $u^{\prime}$ (vertices $u_{i}$ with existing edges $u_{i}\rightarrow u^{\prime}$), and for each viable $u_{i}$, adds new edge $e_{ui}=u\rightarrow u_{i}$ with $\beta(e_{ui})=\beta(e)$. It then tests if the alternative path $u\rightarrow u_{i}\rightarrow u^{\prime}$ achieves the threshold. As illustrated in Figure 2(b), this creates alternative routes to the target using only existing vertices, maintaining all original safety constraints. OrRefine can iteratively explore ancestors of $u_{i}$ if the direct connection fails.

### v2

# arXiv:2512.01047v2[cs.AI] 27 Feb 2026

@@ 1 Introduction
Reinforcement Learning (RL) algorithms have made tremendous strides in recent years Sutton & Barto (2018); Silver et al. (2016); Mnih et al. (2015); Levine et al. (2016). However, most algorithms assume access to a scalar reward function that must be carefully engineered to make environments amenable to RL—a practice known as reward engineering Ibrahim et al. (2024). This creates challenges in applying RL to new environments where useful reward functions are hard to construct. Furthermore, scalar Markovian rewards cannot provide sufficient feedback for certain tasks Abel et al. (2021); Bowling et al. (2023), leading to growing interest in non-Markovian reward functions Li et al. (2017a); Jothimurugan et al. (2021); Alur et al. (2023).
To make non-Markovian rewards tractable, it is standard to represent them via logical specification formulas that capture the intended task. These approaches, known as specification-guided reinforcement learning Aksaray et al. (2016); Li et al. (2017b); Icarte et al. (2018); Jothimurugan et al. (2019; 2021), derive reward functions from logical specifications. However, this creates two challenges: (i) providing specifications granular enough to guide RL algorithms, and (ii) defining accurate labeling functions mapping environment states to specification predicates. Users often create coarse specifications or labeling functions that, while logically correct, provide insufficient guidance for learning.
@@
AutoSpec works with SpectRL specifications; boolean and sequential combinations of reach-avoid tasks Jothimurugan et al. (2019). Any SpectRL specification decomposes into an abstract graph where edges specify reach-avoid tasks Jothimurugan et al. (2021). AutoSpec identifies problematic edges and applies targeted refinements: either modifying the labeling function for regions (Figure 1) or restructuring the graph to add alternative paths. These problematic edges are identified by employing an exploration-guided strategy that utilizes empirical trajectory data to identify edges in the abstract graph whose reach-avoid tasks (i.e. initial, target or unsafe regions) make it hard to learn a good RL policy. For instance, in Figure  1, the initial MID region of the MID-GOAL edge in the abstract graph is under-specified, as it overlaps with a trap state. The trap state is not immediately obvious as there is a path from MID to GOAL. By analyzing explored traces, AutoSpec identifies problematic start states in MID and refines the region to exclude the trap (as shown in Figure 1), thereby refining the logical reach-avoid specification associated to the MID-GOAL edge.
We prove that all refinements maintain soundness, where satisfaction of the refined specification implies satisfaction of the original. AutoSpec integrates with existing SpectRL-compatible algorithms as demonstrated with DiRL Jothimurugan et al. (2021) and LSTS Shukla et al. (2024).
@@
Related work. Recent years have seen substantial progress in solving RL tasks specified via logical specifications Aksaray et al. (2016); Li et al. (2017b); Icarte et al. (2018); Camacho et al. (2019); Giacomo et al. (2019); Hasanbeig et al. (2022; 2019); Hahn et al. (2019); Jothimurugan et al. (2019; 2021); Xu & Topcu (2019). Many works consider different fragments of Linear Temporal Logic (LTL) or their variants for specifying RL tasks. Icarte et al. (2018); Camacho et al. (2019) consider tasks that can be specified using deterministic finite automata (DFA) and solve them by reward machines, which decompose these tasks and translate them into a reward function. The reward function can then be used to train existing RL algorithms. Li et al. (2017b) considers a variant of LTL called TLTL for specifying tasks and propose a method for translating these specifications into continuous reward functions. Hasanbeig et al. (2022; 2019); Hahn et al. (2019) study the translation of tasks specified in LTL into reward functions. Alur et al. (2022) examines the theoretical questions related to the translation of logical specifications into reward functions.
@@
Jothimurugan et al. (2019) defines the specification language SpectRL, a finitary fragment of LTL and provides justification for using this language to define specifications for RL tasks. A compositional method that decomposes SpectRL specifications into an abstract graph and constructs a reward function for each abstract graph edge was proposed in Jothimurugan et al. (2021). The approach by Toro Icarte et al. (2019) focuses on discovering optimal reward structures through environmental exploration and reward analysis. Compositional methods are further explored by Neary et al. (2022), who propose removing unfulfillable subtasks, and Neary et al. (2023), who introduce verification techniques to certify learned policies. Zikelic et al. (2023b) propose Claps, a compositional method for learning neural network policies with formal guarantees on the satisfaction of SpectRL specifications, thus advancing the applicability to safety-critical RL applications by utilizing prior methods for learning reach-avoid policies with formal guarantees Lechner et al. (2022); Zikelic et al. (2023a); Chatterjee et al. (2023).
@@
Recent advancements also include LTL2Action Vaezipoor et al. (2021), which translates LTL specifications into sequences of tasks for RL agents. Other recent approaches, such as  Qiu et al. (2023) and DeepLTL Jackermeier & Abate (2025), leverage goal-conditioned RL and automata-based architectures to solve complex LTL and $\omega$-regular tasks zero-shot or in multi-task settings. Trainify Jin et al. (2022) employs counterexample-guided abstraction and refinement (CEGAR) to iteratively improve policies by addressing failure cases identified through counterexamples. However, these works primarily focus on learning policies for fixed, well-defined specifications. In contrast, AutoSpec studies the problem of automated logical specification refinement towards improving reward functions obtained by translation from coarse logical RL specifications. Thus, our work is complementary to the works on RL from logical specifications and can be integrated into off-the-shelf specification-guided RL algorithms to improve the performance of learned agents.
@@ 2 Preliminaries
MDPs. A Markov Decision Process (MDP) is a tuple $M=(S,A,P,R,\gamma)$, where $S\subseteq\mathbb{R}^{n}$ is the state space, $A\subseteq\mathbb{R}^{m}$ is the action space, $P:S\times A\times S\rightarrow[0,1]$ is the probabilistic transition function, $R$ is the (possibly non-Markovian) reward function, and $\gamma$ is the discount factor. Let $\eta:S\rightarrow[0,1]$ be the initial state distribution. A trajectory $\zeta$ in $M$ is a sequence of states and actions $\zeta=s_{0},a_{0},s_{1},a_{1},\ldots$ where $s_{i}\in S$ and $a_{i}\in A$. We use $\mathcal{Z}$ to denote the set of all trajectories in $M$ and $\mathcal{Z}_{f}$ to denote the set of all finite trajectories in $M$, which are finite prefixes of trajectories ending in states. A (pure) policy $\pi:\mathcal{Z}_{f}\rightarrow A$ assigns an action to each finite trajectory, and a non-Markovian reward $R:\mathcal{Z}_{f}\rightarrow\mathbb{R}$ assigns a reward to a finite trajectory. The MDP $M$ under any policy $\pi$ gives rise to a probability space over the set of all trajectories in the MDP Puterman (1994). We use $\mathbb{P}^{\pi}$ and $\mathbb{E}^{\pi}$ to denote the probability measure and the expectation operator in this probability space, respectively.
Logical specifications for Reinforcement Learning. In this work, we are solving RL tasks defined by logical specifications. Formally, a logical specification (or, simply, a specification) is a boolean function $\phi:\mathcal{Z}\rightarrow\{\textrm{true},\textrm{false}\}$ which specifies whether a trajectory in the MDP satisfies the specification. We write $\zeta\models\phi$ whenever a trajectory $\zeta$ satisfies the specification $\phi$. The objective of a specification-guided RL task is to find a policy $\pi^{*}$ that maximizes the probability of satisfying the given specification $\phi$, i.e. $\pi^{*}\in\textit{argmax}_{\pi}\mathbb{P}^{\pi}[\zeta\models\phi]$. Specification-guided RL algorithms use the specification to create a dense reward that guides the policy search, and therefore outperform algorithms that cannot leverage the specification for learning and instead require manual reward engineering Jothimurugan et al. (2021); Shukla et al. (2024).
SpectRL specification logic. We consider RL tasks specified in the SpectRL specification logic. SpectRL Jothimurugan et al. (2019) is a fragment of Linear Temporal Logic (LTL) which consists of all boolean and sequential combinations of reach-avoid tasks. Formally, a specification in SpectRL is defined in terms of predicates and specification formulas. An atomic predicate is a function $a:S\rightarrow\{\textrm{true},\textrm{false}\}$ which defines a set of states that satisfy the atomic predicate. A predicate is a boolean combination of atomic predicates, i.e. $b:=a\,\mid\,b_{1}\land b_{2}\,\mid\,b_{1}\lor b_{2}$, where $a$ is an atomic predicate and $b_{1}$ and $b_{2}$ are predicates. Specification formulas in SpectRL are defined by the grammar
@@
where $b$ is a predicate and $\phi_{1}$ and $\phi_{2}$ are specification formulas. Intuitively, $\textrm{achieve }b$” requires the agent to reach a state in which the predicate $b$ is satisfied. The clause $\phi_{1}\textrm{ ensuring }b$” requires the agent to satisfy the specification $\phi$ while only visiting states in which the predicate $b$ is satisfied. The clause $\phi_{1};\phi_{2}$” requires the agent to first satisfy specification $\phi_{1}$ and then satisfy specification $\phi_{2}$. The clause $\phi_{1}\textrm{ or }\phi_{2}$” requires satisfaction of at least one of $\phi_{1}$ or $\phi_{2}$. See Jothimurugan et al. (2019) for the formal definition of the semantics of each clause.
Abstract graphs for SpectRL specifications. It was shown in Jothimurugan et al. (2021) that each specification written in the SpectRL specification logic can be translated into an equivalent abstract graph. An abstract graph is a directed acyclic graph (DAG) whose vertices represent sets of MDP states and whose edges are annotated with sets of safe MDP states. Hence, each abstract graph edge defines a reach-avoid specification, where the task is to reach the set of states defined by the target vertex of the edge starting from the set of states defined by the source vertex of the edge, while staying within the set of safe states defined by the edge.
@@
Given a trajectory $\zeta$ in the MDP and an abstract graph $G=(V,E,\beta,s,t)$, we say that $\zeta$ satisfies abstract reachability for $G$ (written $\zeta\models G$) if it gives rise to a path in $G$ that traverses $G$ from $s$ to $t$ and satisfies the reach-avoid specifications of every traversed edge. It was shown in Jothimurugan et al. (2021) that, given any SpectRL specification $\phi$, one can construct an abstract graph $G$ such that $\zeta\models\phi$ if and only if $\zeta\models G$ holds for each trajectory $\zeta$ in the MDP. Hence, solving an RL task for a SpectRL specification reduces to solving an abstract reachability task in the abstract graph $G$.
Problem statement. Given an MDP $M$ and a SpectRL specification $\phi$, our goal is to learn a policy $\pi$ such that the probability $\mathbb{P}^{\pi}[\zeta\models\phi]$ of a trajectory satisfying the specification is maximized.
@@ 3 Automated Refinement of RL Specifications
Overview of AutoSpec. Algorithm 1 shows the complete AutoSpec framework. The algorithm takes as input an MDP $M$, a SpectRL specification $\phi$, a satisfaction threshold $p$, and any specification-guided RL algorithm $\mathcal{A}$. It first translates $\phi$ into an abstract graph $G$ and uses $\mathcal{A}$ to learn policies for the graph edges. For each edge $e$ where $\mathcal{A}$ learned a policy but that policy fails to achieve satisfaction probability $p$, AutoSpec applies four refinement procedures in sequence: SeqRefine, AddRefine, PastRefine, and OrRefine. This ordering reflects increasing levels of structural modification; from local predicate adjustments to graph topology changes. The first refinement that successfully improves performance above threshold $p$ is applied, the graph is updated, and policies are relearned before proceeding to the next edge.
@@
AutoSpec iterates through all edges $e=u\rightarrow u^{\prime}$ of the abstract graph $G$ for which the specification-guided RL algorithm $\mathcal{A}$ has learned a policy but for which the reach-avoid task satisfaction probability is below the provided probability threshold $p$. For each such edge, AutoSpec performs four refinement procedures that focus on different possible reasons for the edge $e=u\rightarrow u^{\prime}$ being challenging for learning a satisfactory policy.
@@ 3.1 Specification Refinement Subprocedures
ReachRefine collects all states along sampled trajectories that successfully reached the goal region $b$. The refined goal region is computed as $b_{r}=b\cap\text{ConvexHull}(\text{reached states})$, effectively excluding unreachable portions of the original target region. AvoidRefine collects states where trajectories entered unsafe regions (complement of $c$). The refined safe region is computed as $c_{r}=c\setminus\text{ConvexHull}(\text{last }k\text{ unsafe trajectory states})$, where $k$ is a hyperparameter specifying how much of the trajectory tail to remove. This removes demonstrated unsafe areas from the safety region.
@@
OrRefine: Exploiting Alternative Paths. The fourth refinement addresses blocked or infeasible direct paths by leveraging the existing graph structure. When the path through edge $e=u\rightarrow u^{\prime}$ cannot be made satisfactory, OrRefine identifies alternative parents of $u^{\prime}$ (vertices $u_{i}$ with existing edges $e_{i}=u_{i}\rightarrow u^{\prime}$), and for each viable $u_{i}$, adds new edge $e_{new}=u\rightarrow u_{i}$ with $\beta(e_{new})=\beta(e)$ and $\beta(e_{i})=\beta(e)\cap\beta(e_{i})$. It then tests if the alternative path $u\rightarrow u_{i}\rightarrow u^{\prime}$ achieves the threshold. As illustrated in Figure 2(b), this creates alternative routes to the target using only existing vertices, maintaining all original safety constraints. OrRefine can iteratively explore ancestors of $u_{i}$ if the direct connection fails.